# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 3.1. Standard clone
$\mathcal{R}(x_{i})=\frac{1}{2e^{\epsilon_{0}}}\mathcal{R}(x_{1}^{0})+\frac{1}{2e^{\epsilon_{0}}}\mathcal{R}(x_{1}^{1})+(1-\frac{1}{e^{\epsilon_{0}}})\text{LO}(x_{i}).$
Let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ denote the shuffling of $\mathcal{R}$. To compute the privacy amplification provided by the shuffle model, we need to compute $D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1}))$ for a given $\epsilon$. The exact computation is computationally complex, so the researchers seek an upper bound for it. A key property is that hockey-stick divergence satisfies the data processing inequality.
@@
$D_{\alpha}\big{(}f(P)||f(Q)\big{)}\leq D_{\alpha}(P||Q).$
If we can find two probability distributions $P_{0}$ and $P_{1}$ along with a post-processing function $f$ such that $f(P_{0})=\mathcal{M}_{S}(X^{0})$ and $f(P_{1})=\mathcal{M}_{S}(X^{1})$, then it follows that $D_{e^{\epsilon}}(P_{0},P_{1})$ is an upper bound for $D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1}))$. We refer to $(P_{0},P_{1})$ as a reduction pair. Different analysis techniques construct different reduction pairs. We first present an intuitive construction of the reduction pair within the standard clone framework, followed by the formal construction.
Definition 0 (Standard Clone Reduction Pair (Intuitive) (Feldman et al., 2022)).
@@
$A_{0}=\begin{cases}0&\text{w.p. }1\\ 1&\text{w.p. }0\\ 2&\text{w.p. }0\end{cases},\ A_{1}=\begin{cases}0&\text{w.p. }0\\ 1&\text{w.p. }1\\ 2&\text{w.p. }0\end{cases},\text{ and }A_{2}=\begin{cases}0&\text{w.p. }\frac{1}{2e^{\epsilon_{0}}}\\ 1&\text{w.p. }\frac{1}{2e^{\epsilon_{0}}}\\ 2&\text{w.p. }1-\frac{1}{e^{\epsilon_{0}}}\end{cases}$
@@
$C\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon_{0}}}),\ A\sim\text{Bin}(C,\frac{1}{2}).$
@@
Algorithm 1 Post-processing function of standard clone (Feldman et al., 2022), $f_{1}$
@@
An additional observation is that if $\mathcal{R}$ is $\epsilon_{0}$-DP, then $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are similar, hence privacy is further amplified (Feldman et al., 2022). The similarity is characterized by the following lemma:
Lemma 2 ((Kairouz et al., 2017)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be an $\epsilon_{0}$-DP local randomizer and $x_{0},x_{1}\in\mathbb{X}$. Then there exists two probability distributions $\mathcal{Q}_{0},\mathcal{Q}_{1}$ such that
$\mathcal{R}(x_{0})=\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{0}+\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}$
@@
$\mathcal{R}(x_{1})=\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{0}+\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}.$
With the help of Lemma 2, (Feldman et al., 2022) gives the following decomposition for generic local randomizers:
$\mathcal{R}(x_{1}^{0})$ $=\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\mathcal{R}(x_{1}^{1})$ $=\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\forall i\in[2,n]:\mathcal{R}(x_{i})$ $=\frac{1}{2e^{\epsilon_{0}}}\mathcal{Q}_{1}^{0}+\frac{1}{2e^{\epsilon_{0}}}\mathcal{Q}_{1}^{1}+(1-\frac{1}{e^{\epsilon_{0}}})\text{LO}(x_{i}).$
@@
Theorem 3 (Standard Clone Reduction (Feldman et al., 2022)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\epsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}^{C},P_{1}^{C}\big{)}$
where $P_{0}^{C},\ P_{1}^{C}$ are defined as below (with “C” denoting “standard Clone”) :
$C\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon_{0}}}),\quad A\sim\text{Bin}(C,\frac{1}{2}),\quad\text{and }\quad\Delta\sim\text{Bern}(\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}).$
$P_{0}^{C}=(A+\Delta,C-A+1-\Delta),\quad P_{1}^{C}=(A+1-\Delta,C-A+\Delta).$
@@
We can construct a post-processing function from $(P_{0}^{C},P_{1}^{C})$ to $\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$, which is similar to Algorithm 1. The only difference is that $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are replaced by $\mathcal{Q}_{1}^{0}$ and $\mathcal{Q}_{1}^{1}$, respectively. ∎
3.2. Privacy Blanket Framework
The decomposition of the standard clone paradigm is based on the projections of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$, using them as reference points and projecting $\mathcal{R}(x_{i})$ onto these bases (Feldman et al., 2022). In contrast, the decomposition provided by the privacy blanket framework first computes the “common part” of all $\mathcal{R}(x)$ (Balle et al., 2019):
@@ 3.2. Privacy Blanket Framework
Here, $\omega$ and $\gamma$ are referred to as the privacy blanket distribution and the total variation similarity of the local randomizer $\mathcal{R}$ (Balle et al., 2019). Each $\mathcal{R}(x)$ can then be decomposed as:
$\mathcal{R}(x)=(1-\gamma)\text{LO}(x)+\gamma\omega,$
where $\text{LO}(x)$ represents the ”left-over” distribution.
@@
The original proof is formulated using the terminology of the “View” of the server. However, we observe that some subsequent works have misinterpreted its meaning (Koskela et al., 2021). To clarify, we restate the privacy blanket technique using the following notation: probability distributions $P_{0}^{B}$ and $P_{1}^{B}$ (with “B” denoting “Blanket”), along with a post-processing function $f^{B}$.
Definition 0 (Privacy Blanket Reduction Pair (Balle et al., 2019) (Restated)).
Let $\boldsymbol{x}_{-1}=(x_{2},x_{3},\dots,x_{n})$ with the inputs from the last $n-1$ users, $\boldsymbol{y}^{a}=(y_{1}^{a},y_{2},\dots,y_{n})$ where $y_{i}\sim R(x_{i})$ is the output of the i-th user, $a$ indicates that the input of the first user is $x_{1}^{a}$, $a\in\{0,1\}$. Let $\boldsymbol{b}=(b_{2},b_{3},\dots,b_{n})$ be binary values indicating which users sample from the privacy blanket distribution. A multiset $Y_{\boldsymbol{b}}^{a}=\mathcal{S}(y_{1}^{a}\cup\{y_{i}|b_{i}=1\})$.
Observe that the distribution of $Y_{\boldsymbol{b}}^{a}$ depends only on $|\boldsymbol{b}|$ rather than $\boldsymbol{b}$, where $|\boldsymbol{b}|$ represents the number of 1 in $\boldsymbol{b}$. We can rewrite it as $Y_{|\boldsymbol{b}|}^{a}=\mathcal{S}({y_{1}}\cup\{y_{i}|y_{i}\sim\omega,i=1,2,\dots,|\boldsymbol{b}|\})$. Then $P_{0}^{B}$ and $P_{1}^{B}$ are defined below:
$P_{0}^{B}=(|\boldsymbol{b}|,Y^{0}_{|\boldsymbol{b}|}),$
$P_{1}^{B}=(|\boldsymbol{b}|,Y^{1}_{|\boldsymbol{b}|}),$
@@
Theorem 5 (Privacy Blanket Reduction (Balle et al., 2019) (Restated)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\epsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}^{B},P_{1}^{B}\big{)}.$
@@
The corresponding post-processing function $f^{B}$ is shown in Algorithm 2. The core idea of the post-processing function $f^{B}$ is that, given $Y_{|\boldsymbol{b}|}^{a}$, it suffices to sample from the left-over distributions of a randomly selected subset of $n-1-|\boldsymbol{b}|$ users and mix the results accordingly. ∎
Algorithm 2 Post-processing function of privacy blanket, $f^{B}$
@@
In Appendix A, we point out and correct the misunderstanding of the privacy blanket framework in (Koskela et al., 2021).
3.3. Vision and failure of stronger clone
The stronger clone is expected to improve the probability of producing a “clone” from $\frac{1}{2e^{\epsilon_{0}}}$ to $\frac{1}{e^{\epsilon_{0}}+1}$. For $\epsilon_{0}>1$, this results in approximately a factor of 2 improvement in the expected number of “clones” (Feldman et al., 2023a). This improvement is anticipated to be achieved through a more refined analysis that, instead of cloning the entire output distributions on differing elements, clones only the portions of those distributions where they actually differ.
Specifically, it leverages a lemma from (Ye and Barg, 2017) to establish the existence of the following decomposition:
Theorem 6 (Corollary 3.4 in (Feldman et al., 2023a)).
Given any $\epsilon_{0}$-DP local randomizer $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and any $n+1$ inputs $x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}\in\mathbb{X}$, if $\mathbb{Y}$ is finite then there exists $p\in[0,1/(e^{\epsilon_{0}}+1)]$ and distributions $\mathcal{Q}_{1}^{0},\mathcal{Q}_{1}^{1},\mathcal{Q}_{1},\mathcal{Q}_{2},\dots,\mathcal{Q}_{n}$ such that
$\mathcal{R}(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\epsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$

### v5

@@ 3.1 Standard clone
$\mathcal{R}(x_{i})=\frac{1}{2e^{\varepsilon_{0}}}\mathcal{R}(x_{1}^{0})+\frac{1}{2e^{\varepsilon_{0}}}\mathcal{R}(x_{1}^{1})+(1-\frac{1}{e^{\varepsilon_{0}}})\text{LO}(x_{i}).$
Let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ denote the shuffling of $\mathcal{R}$. To compute the privacy amplification provided by the shuffle model, we need to compute $D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))$ for a given $\varepsilon$. The exact computation is computationally complex, so the researchers seek an upper bound for it. A key property is that hockey-stick divergence satisfies the data processing inequality.
@@
$\forall\alpha,D_{\alpha}\big{(}f(P)\parallel f(Q)\big{)}\leq D_{\alpha}(P\parallel Q).$
If we can find two probability distributions $P_{0}$ and $P_{1}$ along with a post-processing function $f$ such that $f(P_{0})=\mathcal{M}_{S}(X^{0})$ and $f(P_{1})=\mathcal{M}_{S}(X^{1})$, then it follows that $D_{e^{\varepsilon}}(P_{0}\parallel P_{1})$ is an upper bound for $D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))$. We refer to $(P_{0},P_{1})$ as a reduction pair. Different analysis techniques construct different reduction pairs. We first present an intuitive construction of the reduction pair within the standard clone framework, followed by the formal construction.
Definition 5 (Standard Clone Reduction Pair (Intuitive) [FMT22]).
@@
$A_{0}=\begin{cases}0&\text{w.p. }1\\ 1&\text{w.p. }0\\ 2&\text{w.p. }0\end{cases},\ A_{1}=\begin{cases}0&\text{w.p. }0\\ 1&\text{w.p. }1\\ 2&\text{w.p. }0\end{cases},\text{ and }A_{2}=\begin{cases}0&\text{w.p. }\frac{1}{2e^{\varepsilon_{0}}}\\ 1&\text{w.p. }\frac{1}{2e^{\varepsilon_{0}}}\\ 2&\text{w.p. }1-\frac{1}{e^{\varepsilon_{0}}}\end{cases}$
@@
$C\sim\text{Bin}(n-1,\frac{1}{e^{\varepsilon_{0}}}),\ A\sim\text{Bin}(C,\frac{1}{2}).$
@@
Algorithm 1 Post-processing function of standard clone [FMT22], $f_{1}$
@@
An additional observation is that if $\mathcal{R}$ is $\varepsilon_{0}$-DP, then $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are similar, hence privacy is further amplified [FMT22]. The similarity is characterized by the following lemma:
Lemma 1 ([KOV17]).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be an $\varepsilon_{0}$-DP local randomizer and $x_{0},x_{1}\in\mathbb{X}$. Then there exists two probability distributions $\mathcal{Q}_{0},\mathcal{Q}_{1}$ such that
$\mathcal{R}(x_{0})=\frac{e^{\varepsilon_{0}}}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{0}+\frac{1}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}$
@@
$\mathcal{R}(x_{1})=\frac{1}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{0}+\frac{e^{\varepsilon_{0}}}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}.$
With the help of Lemma 1, [FMT22] gives the following decomposition for generic local randomizers:
$\mathcal{R}(x_{1}^{0})$ $=\frac{e^{\varepsilon_{0}}}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{1}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\mathcal{R}(x_{1}^{1})$ $=\frac{1}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{e^{\varepsilon_{0}}}{e^{\varepsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\forall i\in[2,n]:\mathcal{R}(x_{i})$ $=\frac{1}{2e^{\varepsilon_{0}}}\mathcal{Q}_{1}^{0}+\frac{1}{2e^{\varepsilon_{0}}}\mathcal{Q}_{1}^{1}+(1-\frac{1}{e^{\varepsilon_{0}}})\text{LO}(x_{i}).$
@@
Theorem 1 (Standard Clone Reduction [FMT22]).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\varepsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{C}}\parallel P_{1}^{\mathsf{C}}\big{)}$
where $P_{0}^{\mathsf{C}},\ P_{1}^{\mathsf{C}}$ are defined as below (with “$\mathsf{C}$” denoting “standard Clone”) :
$C\sim\text{Bin}(n-1,\frac{1}{e^{\varepsilon_{0}}}),\quad A\sim\text{Bin}(C,\frac{1}{2}),\quad\text{and }\quad\Delta\sim\text{Bern}(\frac{e^{\varepsilon_{0}}}{e^{\varepsilon_{0}}+1}).$
$P_{0}^{\mathsf{C}}=(A+\Delta,C-A+1-\Delta),\quad P_{1}^{\mathsf{C}}=(A+1-\Delta,C-A+\Delta).$
@@
We can construct a post-processing function from $(P_{0}^{\mathsf{C}},P_{1}^{\mathsf{C}})$ to $\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$, which is similar to Algorithm 1. The only difference is that $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are replaced by $\mathcal{Q}_{1}^{0}$ and $\mathcal{Q}_{1}^{1}$, respectively.
Since the standard clone is a special case of the general clone in Section 4, the correctness of this post-processing function is implied by the proof of Theorem 4. ∎
3.2 Privacy blanket framework
The decomposition of the standard clone is constructed by treating $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ as reference points and projecting each $\mathcal{R}(x_{i})$ onto the basis defined by these two distributions [FMT22]. In contrast, the decomposition provided by the privacy blanket framework takes a different approach: it first identifies the “common part” shared across all $\mathcal{R}(x)$ distributions [BBGN19]:
@@ 3.2 Privacy blanket framework
Here, $\omega$ and $\gamma$ are referred to as the privacy blanket distribution and the total variation similarity of the local randomizer $\mathcal{R}$ [BBGN19]. Each $\mathcal{R}(x)$ can then be decomposed as:
$\mathcal{R}(x)=\gamma\omega+(1-\gamma)\text{LO}(x),$
where $\text{LO}(x)$ represents the “left-over” distribution.
@@
The original proof is formulated using the terminology of the “View” of the server. We restate the privacy blanket technique using the following notation: probability distributions $P_{0}^{\mathsf{B}}$ and $P_{1}^{\mathsf{B}}$ (with “$\mathsf{B}$” denoting “Blanket”), along with a post-processing function $f^{\mathsf{B}}$.
Definition 6 (Privacy Blanket Reduction Pair [BBGN19] (Restated)).
Let $\boldsymbol{x}_{-1}=(x_{2},x_{3},\dots,x_{n})$ with the inputs from the last $n-1$ users, $\boldsymbol{y}^{a}=(y_{1}^{a},y_{2},\dots,y_{n})$ where $y_{i}\sim R(x_{i})$ is the output of the i-th user, $a$ indicates that the input of the first user is $x_{1}^{a}$, $a\in\{0,1\}$. Let $\boldsymbol{b}=(b_{2},b_{3},\dots,b_{n})$ be binary values indicating which users sample from the privacy blanket distribution. A multiset $Y_{\boldsymbol{b}}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|b_{i}=1\})$.
Observe that the distribution of $Y_{\boldsymbol{b}}^{a}$ depends only on $|\boldsymbol{b}|$ rather than $\boldsymbol{b}$, where $|\boldsymbol{b}|$ represents the number of 1 in $\boldsymbol{b}$. We can rewrite it as $Y_{|\boldsymbol{b}|}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|y_{i}\sim\omega,i=1,2,\dots,|\boldsymbol{b}|\})$. Then $P_{0}^{\mathsf{B}}$ and $P_{1}^{\mathsf{B}}$ are defined below:
$P_{0}^{\mathsf{B}}=(|\boldsymbol{b}|,Y^{0}_{|\boldsymbol{b}|}),$
$P_{1}^{\mathsf{B}}=(|\boldsymbol{b}|,Y^{1}_{|\boldsymbol{b}|}),$
@@
Theorem 2 (Privacy Blanket Reduction [BBGN19] (Restated)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\varepsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}}\big{)}.$
@@
The corresponding post-processing function $f^{\mathsf{B}}$ is shown in Algorithm 2. The core idea of the post-processing function $f^{\mathsf{B}}$ is that, given $Y_{|\boldsymbol{b}|}^{a}$, it suffices to sample from the left-over distributions of a randomly selected subset of $n-1-|\boldsymbol{b}|$ users and mix the results accordingly.
In Section 4, we will show that the blanket is, in fact, also a special case of the general clone. The correctness of the corresponding post-processing function $f^{\mathsf{B}}$ follows from the proof of Theorem 4. ∎
Algorithm 2 Post-processing function of privacy blanket, $f^{\mathsf{B}}$
@@
3.3 Vision and failure of stronger clone
The stronger clone is expected to improve the probability of producing a “clone” from $\frac{1}{2e^{\varepsilon_{0}}}$ to $\frac{1}{e^{\varepsilon_{0}}+1}$. For $\varepsilon_{0}>1$, this results in approximately a factor of 2 improvement in the expected number of “clones” [FMT23a]. This improvement is anticipated to be achieved through a more refined analysis that, instead of cloning the entire output distributions on differing elements, clones only the portions of those distributions where they actually differ.
Specifically, it leverages a lemma from [YB17] to establish the existence of the following decomposition:
Theorem 3 (Corollary 3.4 in [FMT23a]).
Given any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and any $n+1$ inputs $x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}\in\mathbb{X}$, if $\mathbb{Y}$ is finite then there exists $p\in[0,1/(e^{\varepsilon_{0}}+1)]$ and distributions $\mathcal{Q}_{1}^{0},\mathcal{Q}_{1}^{1},\mathcal{Q}_{1},\mathcal{Q}_{2},\dots,\mathcal{Q}_{n}$ such that
$\mathcal{R}(x_{1}^{0})$ $=e^{\varepsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\varepsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\varepsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\varepsilon_{0}}p)\mathcal{Q}_{1},$